OpenStudy (anonymous):

find the solution of this system of equations 4x+4y=-16 -8x-6y=-20 juss qive me tha anwser please

OpenStudy (smurfy14):

x=22 y=-26

OpenStudy (anonymous):

Time out.. before we help what do you think you should do.

OpenStudy (anonymous):

smufy let her try these first... please

OpenStudy (anonymous):

well we have to find out what x is first

OpenStudy (anonymous):

riqht

OpenStudy (anonymous):

alright so how would you go about finding x?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

It's not x first or y first, its either of them first,

OpenStudy (anonymous):

Alright in order for you to solve for x you have to get rid of y. Do know the elimination theorum or the substitution theorum yet?

OpenStudy (anonymous):

this is the first step to any equation in math, so practice this alot!

OpenStudy (anonymous):

you would just add or subtract everything to the other side to isolate one variable. choose which one you want to isolate, then you want to add/subtract everything to the other side. then, divide or multiply as required to get x= or y=

OpenStudy (anonymous):

given: 4x+4y=-16-------(1) -8x-6y=-20-------(2) Now, multiply (1) by 2 equation(1) becomes, 8x+8y=-32 -8x-6y=-20 _______________ 2y=-52, y=-26 now, substitute y in eqn1 so, we get 4x+4(-26)=-16 4x-104=-16 4x= 104-16=88 x = 88/4=22 Therefore, x=22, y=-26 plz gimme a medal and become a fan if u like this answer

OpenStudy (anonymous):

krish = celebrity in training

OpenStudy (anonymous):

what they all said haha good job ppls =)

OpenStudy (anonymous):

Your basically getting one variable to cancel out so you can solve for the other, then plugging itback in for the eliminated variable.

OpenStudy (anonymous):

abigail2010 "You're". lol. ;)

OpenStudy (anonymous):

you know what i dont care haha but thanks.. im not an english major for a reason

OpenStudy (anonymous):

haha. i was j/k :p

OpenStudy (anonymous):

lol its fine your right anyways. just had to prove me wrong i see how it is haha

OpenStudy (anonymous):

U haven't got a good answer?

OpenStudy (anonymous):

i have been recommending that she click on good answer for all of her problems. she hasn't chosen a good answer for a lot of her problems and it's cluttering up the unanswered questions section... i guess everyone doesn't understand this as we do :?

OpenStudy (anonymous):

she has to do it for god sake

OpenStudy (anonymous):

4x+4y=-16 divide by 4 put x on other side y=-x-4 -8x-6y=-20 divide by 2 insert y -4x-3(-x-4)=-10 -4x+3x+12=-10 -x=-22 strike the minus since on both sides, insert into y(x) y=-22-4=-26

OpenStudy (anonymous):

nope. don't see it happening.... :/